OpenStudy (anonymous):

1.Simplify the expression by combining like terms. x - 3y + x + 3y A. 2x + 6y B. 6y C. 2x D. x + 6y 2.Simplify the expression by combining like terms. x - 23y + x + 23y A. 2x + 46y B. x + 46y C. 46y D. 2x 3.Combining like terms to simplify 3y + 2x + 4 + y – x equals 4y + 3x + 4 A. True B. False 4.Decide whether the given number is a solution of the given equation. Is 1 a solution of 4 - p = 3 ? A. Yes B. No

OpenStudy (anonymous):

5.Decide whether the given number is a solution of the given equation. Is 4 a solution of 6x = 28 - x? A. Yes B. No 6.17 is a solution of x + 4 = 21 A. yes B. no 7.Solve the equation. y + 13 = 10 A. y = 3 B. y = -23 C. y = 23 D. y = -3 8. PREALG4M 3.2.2-2+ Solve the equation. 9 = x - 3 A. x = -6 B. x = 12 C. x = 6 D. x = -12 Points Possible:2.00 9. Y = 2 is a solution to the equation 3y = 12 A. True B. False